# MotionDreamer: Zero-Shot 3D Mesh Animation from Video Diffusion Models

### v1

@@ Abstract
Animation techniques bring digital 3D worlds and characters to life. However, manual animation is tedious and automated techniques are often specialized to narrow shape classes. In our work, we propose a technique for automatic re-animation of arbitrary 3D shapes based on a motion prior extracted from a video diffusion model. Unlike existing 4D generation methods, we focus solely on the motion, and we leverage an explicit mesh-based representation compatible with existing computer-graphics pipelines. Furthermore, our utilization of diffusion features enhances accuracy of our motion fitting. We analyze efficacy of these features for animation fitting and we experimentally validate our approach for two different diffusion models and four animation models. Finally, we demonstrate that our time-efficient zero-shot method achieves a superior performance re-animating a diverse set of 3D shapes when compared to existing techniques in a user study. The project website is located at https://lukas.uzolas.com/MotionDreamer.
@@ 1 Introduction
We build on the remarkable success of Diffusion models [7, 8]. Beyond producing nearly photo-realistic 2D images [9, 10, 11, 12], diffusion was also adapted for 3D [13, 14] and 4D shape synthesis [15, 16, 17, 18, 19, 20, 21, 22, 23]. However, the associated methods suffer from either a high optimization cost and low diversity [24] of the mode-seeking Stochastic Distillation Sampling [13] (SDS), or, as we show in our paper, they are susceptible to the visual artifacts in RGB outputs of existing VDMs.
Instead, we leverage the surprising versatility of semantic features extracted from diffusion models for down-stream tasks such as zero-shot segmentation or semantic feature matching [25], which we adapt for motion fitting. We rely on a classical surface mesh representation in combination with diverse animation models [26, 27, 28, 29] to obtain animated 3D shapes that are fast to render, compatible with existing rendering frameworks and versatile across object classes.
@@ 2.2 Shape and pose representations
While different poses and motion can be encoded as a sequence of static representations [46], specialized representations ease editing for both arbitrary and class-specific shapes. In the first category, deformation fields offer maximal flexibility for dense volumetric optimization [47], Neural Jacobian Fields (NJFs) [26] offer space-time continuity and smoothness for surface optimization and external cages reduce the control space for easier editing [48]. In the second category, low-dimensional templates support manual animation and motion capture by combining Linear Blend Skinning [49] and Blend Shapes [50, 51] for narrow classes of shapes such as faces [52, 29], bodies [27, 53], hands [54], or even animals [28]. Our method is agnostic to the choice of an animation model, which we test on high-dimensional NJFs [26] and on low-dimensional templates [27, 28, 29].
@@
Learned priors can also be used for text-conditioned motion synthesis [75]. However, this is in practice limited to human domain [76, 77, 78, 79] where annotated 3D motion datasets exist [80, 81] or to other skeletal shapes [82] if at least 2D annotations are available. Alternatively, image and video generative models enabled class-agnostic joint shape and motion 4D generation [15, 16, 17, 18, 19, 20, 21, 22, 23] usually based on Stochastic Distillation Sampling (SDS) [13] which, however, narrows the sampled distribution [24]. Closest to us, Ren et al. [17] extract motion from a full video input. Our method shares the idea of extracting motion from a video model but we show better performance thanks to utilizing feature space. Additionally, by focusing on motion alone we achieve faster sampling.
Finally, both captured or generated motion can be transferred from one shape to another [83], either based on morphological similarity [84, 85] or data-driven domain matching [86, 87]. We experimentally show that our method is preferable when neither of the two conditions can be satisfied.
@@ 4 Method
We define the input mesh $\mathcal{M}$ as a tuple of $N$ vertices and $M$ triangular faces :=($\mathcal{M}{}:=(\{\mathbf{u}_{n}\in\mathbb{R}^{3}|n=0,...,N-1\},\{\mathbf{f}_{m}\in\mathbb{N}^{3}|m=0,...,M-1\})$. Next, we define $\tau:(\mathcal{M}{},\mathbf{p}{})\rightarrow\mathcal{M}^{\prime}$ as a function transforming vertices to produce a mesh :=($\mathcal{M}^{\prime}:=(\{\mathbf{u}_{n}^{\prime}\},\{\mathbf{f}_{m}\})$ with a novel pose described by animation parameters $\mathbf{p}{}\in\mathbb{R}^{P}$. We refer to $\mathbf{p}_{\textrm{init}}$ as the input pose where $\tau(\mathcal{M},\mathbf{p}_{\textrm{init}})\equiv\mathcal{M}$ and, without a loss of generality, we assume it matches the first frame. Finally, $r_{rgb}:(\mathcal{M}{},\mathcal{C}{},\mathcal{T},\mathbf{B})\rightarrow\mathbf{I}{}_{rgb}{}$ is a rendering function producing an RGB image $\mathbf{I}{}_{rgb}{}\in\mathbb{R}^{H\times W\times 3}$ of the mesh $\mathcal{M}{}$ for a camera $\mathcal{C}{}$, surface texture $\mathcal{T}$, and a background image $\mathbf{B}\in\mathbb{R}^{H\times W\times 3}$.
@@ 4.2 Motion Generation
The motion produced by our method originates from a VDM conditioned by our rendered mesh image $\mathbf{I}{}_{rgb}^{0}=r_{rgb}(\mathcal{M}{}^{0},\mathcal{C}{},\mathcal{T},\mathbf{B})$ and an embedding of the intended motion text description. We sample the generator in a multi-step diffusion process over $T$ step denoted as $t\in[0,...,T-1]$ with scheduling details specific to each VDM. Because the temporally incoherent visual artifacts in RGB video outputs make motion tracking difficult (see Fig. 1), we extract semantically meaningful U-Net features $\mathbf{A}^{t}_{u}$ at time step $t=\hat{t}$ and U-Net layer $u=\hat{u}$ as explained in Sec. 3.2, and we show that this improves the fitting accuracy. We motivate our choice of $\hat{t}$ and $\hat{u}$ in Sec. 5.3, and will omit the suffixes from now on for brevity, such that $\hat{\mathbf{A}}\in\mathbb{R}^{L\times\hat{H}\times\hat{W}\times\hat{A}}\equiv\mathbf{A}^{\hat{t}}_{\hat{u}}$ and $\hat{\mathbf{A}}^{l}$ selects the video frame $l$ of $L$. We further assume $\hat{\mathbf{A}}^{0}$ corresponds to the input image $\mathbf{I}{}_{rgb}^{0}$ (see Appendix A.1 for a discussion).
@@ 4.3 Motion fitting
Our pose fitting is based on reprojection of $\hat{\mathbf{A}}^{0}$ to any new pose $\mathbf{p}^{l}$. First, we use projective texturing to map $\hat{\mathbf{A}}^{0}$ to $\mathcal{M}$. We obtain per-vertex features $\{\mathbf{a}_{n}\}$ by mapping each mesh vertex $\mathbf{u}_{n}$ to the image plane of the camera $\mathcal{C}$ and sampling $\hat{\mathbf{A}}^{0}$ as $\mathbf{a}_{n}=\hat{\mathbf{A}}^{0}[P(\mathbf{u}_{n},\mathcal{C})],$ where $P(.)$ is a world space to image plane projection function and [.][.][ . ] is a bilinear sampler. Finally, we transform $\mathcal{M}$ to $\mathcal{M}^{l}=\tau(\mathcal{M},\mathbf{p}^{l})$ for a given novel pose $\mathbf{p}^{l}$ and we render a feature image
@@
Pose Optimization
We observe that direct optimization of each $\mathbf{p}{}^{l}$ independently is prone to local minima. Instead, we exploit the implicit bias of Multi-Layer-Perceptrons (MLPs) towards smooth functions, and regress $\mathbf{p}{}^{l}$ as a frame-dependent offset from an initial pose $\mathbf{p}_{\textrm{init}}$ such that $\mathbf{p}{}^{l}=\alpha m_{\omega}(\gamma(l))+\mathbf{p}_{\textrm{init}}$, where $\alpha=0.01$ is a scaling constant, $\gamma$ is a frequency encoding [42], and $m(.)$ is an MLP with learnable parameters $\omega$. We optimize $\omega$ by gradient descent to enforce semantic correspondences $\mathbf{I}{}_{feat}^{l}\approx\hat{\mathbf{A}}^{l}\textit{ }:\textit{ }$ between the animated mesh and the video using the Rendering loss
@@ 4.4 Implementations Details
We evaluate 2 VDMs: VideoComposer [3] (VC) and DynamiCrafter [4] (DC) with $\hat{\mathbf{A}}$ resolution of $(160,88)$ and $(128,72)$ respectively ($1/8$ of their outputs). We use their recommended inference schedulers with $T=50$ steps. Our assumption of $\hat{\mathbf{A}}^{0}\sim\mathbf{I}{}_{rgb}^{0}$ is satisfied by design for VC, and we present a solution for DC in Appendix A.1. We empirically find VC performs better for images with the background $\mathbf{B}$, while DC performs well even with a uniform white background (see Appendix D.1).
@@ 5.1 User Study
We compare our method to two other approaches for zero-shot 3D motion synthesis. First, we compare to DG4D [17] as an end-to-end shape-and-motion generative method based on image and video diffusion. Second, in absence of a class-agnotic method, we compare to a human motion diffusion model (MDM) [79] combined with motion retargeting (MT) [84]. We run our method with both VC and DC backbones and use the same generated videos as inputs for DG4D (see Appendix B.1 for details). We use 9 meshes and a total of 12 prompts combined to obtain 2 human stimuli (using the SMPL mesh), 2 horses (SMAL), 2 faces (FLAME) and 4 other stimuli each with a unique mesh (NJF). See Fig. 1 and Fig. 3 for visual examples and Appendix B.2 for a complete list.
@@ Results
As summarized in Tbl. 5.2, Ours consistently achieves better results with textured inputs than with untextured inputs, which motivates our Single-View Texturing (Sec. 4.1). Furthermore, Ours (full) with semantic features $\hat{\mathbf{A}}$ achieves lower errors than the variant with RGB features, which documents the utility of these features for our task. Finally, Ours (full) compares favorably even to the WHAM pose estimator despite the lack of human-specific training. This might be explained by the artificial appearance of our input videos which differ from common human pose estimation datasets. We do not claim general supremacy of our method for human pose estimation. This is showcased in Fig. 5 (right), where our method struggles to avoid physiologically implausible poses.
@@
[…]
figure Left: The pose fitting performance of WHAM [67] and variants of our method for re-rendered AIST++ human body sequences [95]. Less is better for all metrics (see Sec. 5.2). Right: PA-MPJPE $\downarrow$ with a standard deviation range for features $\mathbf{A}^{t}_{\hat{u}}$ extracted for different diffusion steps $t$.
@@
[…]
figure Left: Performance of our ablated method variants in pose fitting. Notable performance impacts highlighted in red. Right: Depth regularization prevents undesirable motion-in-depth explanations.
@@ 5.3 Ablations
We reuse the pose optimization experiment to validate our design choices. To this end, we follow the same procedure for 6 of the same AIST++ sequences [95]. First, we analyze the choice of $\hat{u}$ (Fig. 4 right) and $\hat{t}$ (Fig. 5.2 right) for extraction of $\hat{\mathbf{A}}$ using PA-MPJPE. We observe the best performance for $\hat{u}=3$, which we consequently use for both VDMs in our other experiments. We further find our method is not sensitive to the choice of $\hat{t}$ above $t\approx 15$. Therefore, we select $\hat{t}=20$ for VC and $\hat{t}=40$ for DC.
Next, we ablate our regularization losses (Fig. 5.2 left). As expected, the smoothness of $\mathcal{L}_{s}$ reduces the Acceleration error, while $\mathcal{L}_{f}$ reduces shape distortions recorded by the Per-Vertex Error. In contrast, the depth regularization of $\mathcal{L}_{d}$ does not lead to an improvement in performance metrics, but we observe that it prevents perceptually-objectionable depth errors (Fig. 5.2 right).
@@ 6 Discussion
Single-view motion supervision struggles to resolve motion-in-depth or occlusions, which we mitigate using regularization at a risk of overall motion reduction (see Fig. 5.2 right). Alternatively, a multi-view supervision [82] could offer an improvement at the cost of additional training data. We demonstrate a zero-shot method supporting a range of animation models, but we acknowledge that the high degree-of-freedom in NJF permits undesired distortions (see Fig. 5a). These could be potentially remedied though a static shape supervision inspired by 3D generative models [14] with a possible diversity reduction stemming from SDS [24]. On top of this, the motion produced by the current VDMs might not adhere to the prompt or might contain physically impossible transitions (see Fig. 5b). To counter this, the fast run-time of our method could be combined with a suitable rejection heuristic. Furthermore, we expect to benefit from future VDM improvements [39, 40]. This will also allow for generating longer sequences, necessitating memory off-loading, which is currently absent in our implementation. Finally, an interesting future direction is to close the loop and constrain the VDM generation with our simultaneously optimized 3D animation model in order to prevent any spatiotemporal distortions from emerging.

### v2

# MotionDreamer: Exploring Semantic Video Diffusion features for Zero-Shot 3D Mesh Animation

@@ Abstract
Animation techniques bring digital 3D worlds and characters to life. However, manual animation is tedious and automated techniques are often specialized to narrow shape classes. In our work, we propose a technique for automatic re-animation of various 3D shapes based on a motion prior extracted from a video diffusion model. Unlike existing 4D generation methods, we focus solely on the motion, and we leverage an explicit mesh-based representation compatible with existing computer-graphics pipelines. Furthermore, our utilization of diffusion features enhances accuracy of our motion fitting. We analyze efficacy of these features for animation fitting and we experimentally validate our approach for two different diffusion models and four animation models. Finally, we demonstrate that our time-efficient zero-shot method achieves a superior performance re-animating a diverse set of 3D shapes when compared to existing techniques in a user study.
@@ 1 Introduction
We build on the remarkable success of Diffusion models [10, 11]. Beyond producing nearly photo-realistic 2D images [12, 13, 14, 15], diffusion was also adapted for 3D [16, 17] and 4D shape synthesis [18, 19, 20, 21, 22, 23, 24, 25, 26]. However, the associated methods suffer from either a high optimization cost and low diversity [27] of the mode-seeking Stochastic Distillation Sampling [16] (SDS), or, as we show in our paper, they are susceptible to the visual artifacts in RGB outputs of existing VDMs. Furthermore, our method generates a unique animation as a sequence of object poses in a matter of minutes rather than hours common for end-to-end 4D generative methods. This is a feature crucial for processing of larger sample sets with subsequent filtering based on subjective preferences. Therefore, we position our approach into a category distinct from end-to-end 4D generation.
Instead of iterative SDS, we leverage the surprising versatility of semantic features extracted from diffusion models for down-stream tasks such as one-shot segmentation [28] or semantic feature matching [29, 30], which we adapt for motion fitting. We rely on a classical surface mesh representation in combination with diverse animation models [2, 31, 3, 1] to obtain animated 3D shapes that are fast to render, compatible with existing rendering frameworks and versatile across object classes.
@@ 2.2 Shape and pose representations
While animation of object poses can be encoded as a sequence of static representations [47], specialized representations ease editing for both arbitrary and class-specific shapes. In the first category, deformation fields offer maximal flexibility for dense volumetric optimization [48], Neural Jacobian Fields (NJFs) [2] offer space-time continuity and smoothness for surface optimization and external cages reduce the control space for easier editing [49]. In the second category, low-dimensional templates support manual animation and motion capture by combining Linear Blend Skinning [50] and Blend Shapes [51, 52] for specific classes of shapes such as faces [53, 1], bodies [31, 54], hands [55], or even animals [3]. Our method is agnostic to the choice of an animation model, which we test on high-dimensional NJFs [2] and on low-dimensional templates [31, 3, 1].
@@
Learned priors can also be used for text-conditioned motion synthesis [76]. However, this is in practice limited to human domain [77, 78, 79, 80] where annotated 3D motion datasets exist [81, 82] or to other skeletal shapes [83] if at least 2D annotations are available. Alternatively, image and video generative models enabled class-agnostic joint shape and motion 4D generation [18, 19, 20, 21, 22, 23, 24, 25, 26] is usually based on Stochastic Distillation Sampling (SDS) [16] which, however, narrows the sampled distribution [27] due it its mode-seeking behavior. Closest to us, Ren et al. [20] extract motion from a full video input. Our method shares the idea of extracting motion from a video model but thanks to utilizing the feature space it produces more natural motion with fewer visual artifacts. Furthermore, we do not use 3D uplifting methods requiring background masks such as Zero-1-to-3 [84]. Additionally, by focusing on motion alone we achieve faster sampling. Finally, both captured or generated motion can be transferred from one shape to another [85], either based on morphological similarity [86, 87] or data-driven domain matching [88, 89]. We experimentally show that our method is preferable when neither of the two conditions can be satisfied.
@@ 4 Method
We define the input mesh $\mathcal{M}$ as a tuple of $N$ vertices and $M$ triangular faces :=($\mathcal{M}{}:=(\{\mathbf{u}_{n}\in\mathbb{R}^{3}|n=0,...,N-1\},\{\mathbf{f}_{m}\in\mathbb{N}^{3}|m=0,...,M-1\})$. Next, we define $\tau:(\mathcal{M}{},\mathbf{p}{})\rightarrow\mathcal{M}^{\prime}$ as a function transforming vertices to produce a mesh :=($\mathcal{M}^{\prime}:=(\{\mathbf{u}_{n}^{\prime}\},\{\mathbf{f}_{m}\})$ with a novel pose described by animation parameters $\mathbf{p}{}\in\mathbb{R}^{P}$. We refer to $\mathbf{p}_{\textrm{init}}$ as the input pose where $\tau(\mathcal{M},\mathbf{p}_{\textrm{init}})\equiv\mathcal{M}$ and, without a loss of generality, we assume it matches the first frame. Finally, $r_{rgb}:(\mathcal{M}{},\mathcal{C}{},\mathcal{T},\mathbf{B})\rightarrow\mathbf{I}{}_{rgb}{}$ is a rendering function producing an RGB image $\mathbf{I}{}_{rgb}{}\in\mathbb{R}^{H\times W\times 3}$ of the mesh $\mathcal{M}{}$ for a manually defined canonical camera $\mathcal{C}{}$, surface texture $\mathcal{T}$, and a background image $\mathbf{B}\in\mathbb{R}^{H\times W\times 3}$.
@@ 4.2 Motion Generation
The motion produced by our method originates from a VDM conditioned by our rendered mesh image $\mathbf{I}{}_{rgb}^{0}=r_{rgb}(\mathcal{M}{}^{0},\mathcal{C}{},\mathcal{T},\mathbf{B})$ and an embedding of the intended motion text description. We sample the generator in a multi-step diffusion process over $T$ steps denoted as $t\in[0,...,T-1]$ with scheduling details specific to each VDM. Because the temporally incoherent visual artifacts in RGB video outputs make motion tracking difficult (see Fig. 1), we extract semantically meaningful U-Net features $\mathbf{A}^{t}_{u}$ at time step $t=\hat{t}$ and U-Net layer $u=\hat{u}$ as explained in Sec. 3.2, and we show that this improves the fitting accuracy. We motivate our choice of $\hat{t}$ and $\hat{u}$ in Sec. 5.3, and will omit the suffixes from now on for brevity, such that $\hat{\mathbf{A}}\in\mathbb{R}^{L\times\hat{H}\times\hat{W}\times\hat{A}}\equiv\mathbf{A}^{\hat{t}}_{\hat{u}}$ and $\hat{\mathbf{A}}^{l}$ selects the video frame $l$ of $L$. We further assume $\hat{\mathbf{A}}^{0}$ corresponds to the input image $\mathbf{I}{}_{rgb}^{0}$ (see Appendix A.1 for a discussion).
@@ 4.3 Motion fitting
Our mesh pose fitting is based on reprojection of $\hat{\mathbf{A}}^{0}$ to any new pose $\mathbf{p}^{l}$. First, we use projective texturing to map $\hat{\mathbf{A}}^{0}$ to $\mathcal{M}$. We obtain per-vertex features $\{\mathbf{a}_{n}\}$ by mapping each mesh vertex $\mathbf{u}_{n}$ to the image plane of the camera $\mathcal{C}$ and sampling $\hat{\mathbf{A}}^{0}$ as $\mathbf{a}_{n}=\hat{\mathbf{A}}^{0}[P(\mathbf{u}_{n},\mathcal{C})],$ where $P(.)$ is a world space to image plane projection function and [.][.][ . ] is a bilinear sampler. Finally, we transform $\mathcal{M}$ to $\mathcal{M}^{l}=\tau(\mathcal{M},\mathbf{p}^{l})$ for a given novel pose $\mathbf{p}^{l}$ and we render a feature image
@@
Mesh Pose Optimization
We observe that direct optimization of each $\mathbf{p}{}^{l}$ independently is prone to local minima. Instead, we exploit the implicit bias of Multi-Layer-Perceptrons (MLPs) towards smooth functions, and regress $\mathbf{p}{}^{l}$ as a frame-dependent offset from an initial pose $\mathbf{p}_{\textrm{init}}$ such that $\mathbf{p}{}^{l}=\alpha m_{\omega}(\gamma(l))+\mathbf{p}_{\textrm{init}}$, where $\alpha=0.01$ is a scaling constant, $\gamma$ is a frequency encoding [43], and $m(.)$ is an MLP with learnable parameters $\omega$. We optimize $\omega$ by gradient descent to enforce semantic correspondences between the animated mesh and the video, i.e. $\mathbf{I}{}_{feat}^{l}\approx\hat{\mathbf{A}}^{l}$, using the rendering loss:
@@ 4.4 Implementations Details
We evaluate 2 VDMs: VideoComposer [6] (VC) and DynamiCrafter [7] (DC) with $\hat{\mathbf{A}}$ resolution of $(160,88)$ and $(128,72)$ respectively ($1/8$ of their outputs). We use their recommended inference schedulers with $T=50$ steps. Our assumption of $\hat{\mathbf{A}}^{0}\sim\mathbf{I}{}_{rgb}^{0}$ is satisfied by design for VC, and we present a solution for DC in Appendix A.1. We empirically find VC performs better for images with the background $\mathbf{B}$, while DC performs well even with a uniform white background. Additionally, we assessed another VDM, Stable Video Diffusion [5], but we discarded due to its low motion quality (see Appendix D.1, D.5).
@@ 5.1 User Study
We compare our method to two other approaches for zero-shot 3D motion synthesis. First, we compare to DG4D [20] as an end-to-end shape-and-motion generative method based on image and video diffusion. Second, in absence of a class-agnotic method, we compare to a human motion diffusion model (MDM) [80] combined with motion retargeting (MT) [86]. We provide an additional qualitative comparison to a contemporary end-to-end generative method Consistent4D [25] in Appendix D.2. We run our method with both VC and DC backbones and use the same generated videos as inputs for DG4D (see Appendix B.1 for details). We use 9 meshes and a total of 12 prompts combined to obtain 2 human stimuli (using the SMPL mesh), 2 horses (SMAL), 2 faces (FLAME) and 4 other stimuli each with a unique mesh (NJF). See Fig. 1, Fig. 3, and the supplemental video for visual examples, Appendix B.2 for a complete list.
@@ Results
As summarized in Tbl. 1, Ours consistently achieves better results with textured inputs than with untextured inputs, which motivates our Single-View Texturing (Sec. 4.1). Furthermore, Ours (full) with semantic features $\hat{\mathbf{A}}$ achieves lower errors than the variant with RGB features, which documents the utility of these features for our task. Moreover, Ours (full) compares favorably even to the WHAM pose estimator despite the lack of human-specific training. This might be explained by the artificial appearance of our input videos which differ from common human pose estimation datasets. We do not claim general supremacy of our method for human pose estimation. This is showcased in Fig. 8 (right), where our method struggles to avoid physiologically implausible poses. Finally, in Fig. 6 we compare both features qualitatively in our full generative method and confirm that our semantic features lead to a better motion fit with fewer artifacts. See Appendix D.3 for more examples.
@@
Table 1: The pose fitting performance of WHAM [68] and variants of our method for re-rendered AIST++ human body sequences [97]. Less is better for all metrics (see Sec. 5.2).
Figure 5: Left: PA-MPJPE $\downarrow$ with a standard deviation range for features $\mathbf{A}^{t}_{\hat{u}}$ extracted for different diffusion steps $t$. Right: Depth regularization prevents undesirable motion-in-depth explanations.
@@
Table 2: Performance of our ablated method variants in pose fitting. Notable performance impacts highlighted in red.
@@ 5.3 Ablations
Figure 6: Effect of features on optimization under different guidance scales. Higher guidance scales are usually accompanied by stronger artifacts in RGB space (right column). The top row shows frame 11 of the VDM output, i.e. the target expression. The consecutive rows show semantic features vs RGB used for optimization.
We reuse the pose optimization experiment to validate our design choices. To this end, we follow the same procedure for 6 of the same AIST++ sequences [97]. First, we analyze the choice of $\hat{u}$ (Fig. 4 right) and $\hat{t}$ (Fig. 5 left) for extraction of $\hat{\mathbf{A}}$ using PA-MPJPE. We observe the best performance for $\hat{u}=3$, which we consequently use for both VDMs in our other experiments. We further find our method is not sensitive to the choice of $\hat{t}$ above $t\approx 15$. Therefore, we select $\hat{t}=20$ for VC and $\hat{t}=40$ for DC.
Next, we ablate our regularization losses (Tbl. 2). As expected, the smoothness of $\mathcal{L}_{s}$ reduces the Acceleration error, while $\mathcal{L}_{f}$ reduces shape distortions recorded by the Per-Vertex Error. In contrast, the depth regularization of $\mathcal{L}_{d}$ does not lead to an improvement in performance metrics, but we observe that it discourages perceptually-objectionable depth errors (Fig. 5 right).
Finally, in Fig. 8 we ablate the mesh resolution in our method with NJF. We find that the output quality degrades gracefully and predictably with reduced vertex count which is favorable for our method’s versatility. See Appendix D.4 for an extended discussion.
@@ 6 Discussion
Figure 7: Effect of number of vertices on our method. Model source: Jaka Ardian 3D art / model from Indonesia.
@@
Single-view motion supervision struggles to resolve motion-in-depth or occlusions, which we mitigate using regularization at a risk of overall motion reduction (see Fig. 5 right). We acknowledge this as a limitation and a motivation for further research which could offer an improvement through multi-view supervision at the cost of additional training data [83]. We demonstrate a zero-shot method supporting a range of animation models, but we acknowledge that the high degree-of-freedom in NJF permits undesired distortions (see Fig. 8a). These could be potentially remedied though a static shape supervision inspired by 3D generative models [17] with a possible diversity reduction stemming from SDS [27]. On top of this, the motion produced by the current VDMs might not adhere to the prompt or might contain physically impossible transitions (see Fig. 8b). To counter this, the fast run-time of our method could be combined with a suitable rejection heuristic. Furthermore, we expect to benefit from future VDM improvements [40, 41]. This will also allow for generating longer sequences, necessitating memory off-loading, which is currently absent in our implementation. Finally, an interesting future direction is to close the loop and constrain the VDM generation with our simultaneously optimized 3D animation model in order to prevent any spatiotemporal distortions from emerging.